### v1

# From Prompt Engineering to Prompt Science With Human in the Loop

Abstract.
As LLMs make their way into many aspects of our lives, one place that warrants increased scrutiny with LLM usage is scientific research. Using LLMs for generating or analyzing data for research purposes is gaining popularity. But when such application is marred with ad-hoc decisions and engineering solutions, we need to be concerned about how it may affect that research, its findings, or any future works based on that research. We need a more scientific approach to using LLMs in our research. While there are several active efforts to support more systematic construction of prompts, they are often focused more on achieving desirable outcomes rather than producing replicable and generalizable knowledge with sufficient transparency, objectivity, or rigor. This article presents a new methodology inspired by codebook construction through qualitative methods to address that. Using humans in the loop and a multi-phase verification processes, this methodology lays a foundation for more systematic, objective, and trustworthy way of applying LLMs for analyzing data. Specifically, we show how a set of researchers can work through a rigorous process of labeling, deliberating, and documenting to remove subjectivity and bring transparency and replicability to prompt generation process. A set of experiments are presented to show how this methodology can be put in practice.
Large Language Models (LLMs), Prompt engineering, Qualitative coding, Verification
copyright: rightsretainedpublicationmonth: Decemberjournalyear: 2023ccs: Computing methodologies Information extractionccs: Human-centered computing User centered design
1. Introduction
Large Language Models (LLMs), in the recent years, have become more sophisticated and capable for them to be applicable in many situations and tasks. These tasks are not limited to information extraction and synthesis (Lindemann, 2023), but also expanded to analysis, creation, and reasoning (Chew et al., 2023). Unsurprisingly, many researchers have found their uses for various research tasks. LLMs are being used in identifying relevant papers (Paroiu et al., 2023), synthesizing literature reviews (Antu et al., 2023), writing proposals (Gómez-Rodríguez and Williams, 2023), and analyzing data (Shen et al., 2023). They have been found effective and useful in investigative tasks such as drug discovery (Vert, 2023; Savage, 2023), and while such capabilities and wide applicabilities of LLMs have opened up new avenues for supporting research, there is a growing concern that a large portion of this success hinges on prompt engineering, which is often an ad-hoc method to revise prompts being fed into an LLM to achieve desired response or analysis (Reynolds and McDonell, 2021).
Using such a process for scientific research could be dangerous. At best, there is a possibility of creating a feedback loop with self-fulfilling prophesy. At worst, one may be overfitting hypotheses to data, leading to untrustworthy claims and findings. In most cases, overreliant on prompt engieering could lead to unexplainable, unverifiable, and less generalizable outcomes (Patel and Sattler, 2023; Shah and Bender, 2023), lacking a scientific approach to research.
Here, we consider a scientific approach to be the one that is well-documented and transparent, verifiable, repeatable, and free of individual biases and subjectivity. We argue that most prompt engineering techniques do not lead to a scientific approach to research as they violate at least one of these principles. While such ad-hocness can be bad in most situation, when it comes to scientific research that expects a certain level of vigor, it is especially problematic.
If we want to continue harnessing the powers of LLMs, we need to do this in a responsible manner with enough scientific rigor. We focus on situations where data is fed into an LLM with a set of instructions or a prompt to label or analyze that data. Building on an LLM’s ability to summarize or synthesize data, many recent research projects have deployed them for such tasks (Reynolds and McDonell, 2021; Shen et al., 2023). There are also efforts to use LLMs for generating datasets and benchmarks (Thomas et al., 2023; Peng et al., 2023). These could be foundations for future research. Therefore, it is paramount that such outcomes from LLMs are thoroughly validated and proven trustworthy.
In this article, we demonstrate how to do this by turning an ad-hoc method of prompt engineering to verifiable and replicable method of prompt science. Specifically, we take inspiration from a rich literature and methods of qualitative coding to employ human-in-the-loop for verifying (1) responses from LLMs; and (2) systematically constructing prompts.
2. Prompt engineering prospects and problems
LLMs are built on transformer architecture that contain encoders and decoders modules. They are trained on large amounts of text corpora. A lot of what an LLM generates heavily depends on the inner workings of this architecture and the data used for training. Most users and researchers do not have the ability to directly change that architecture or the training process. However, they can apply various techniques such as fine-tuning (Hu et al., 2023; Ding et al., 2023) and prompt engineering (Strobelt et al., 2022) to shape the outputs as they desire.
Figure 1. A typical flow of prompt engineering for research process that involves labeling data.
Figure 1 shows a conceptual flow of a typical prompt revision or engineering in a research setting. Here, a researcher is interested in getting some input data analyzed or labeled for a downstream application. To get the desired outcomes, the set of instructions or the prompt provided to the LLM is interatively revised. This has been used by many recent works such as (Shah et al., 2023a; Ziems et al., 2023; Chew et al., 2023; Ma et al., 2023). While this could accomplish the task, it leaves us with many potential issues. For instance, if the prompt is revised by a single researcher in ad-hoc manner just to fine-tune the output, there may be personal biases and subjectivity introduced. In addition, if this process is not well-documented, it may be hard to understand or mitigate these issues. This may also make it harder for someone to verify or replicate the process. These issues may hinder scientific progress that relies on systematic knowledge being generated, questioned, and shared. There may also be issues of self-fulfilling prophecy or feedback loop that are especially dangerous when we are dealing with highly opaque systems such as an LLM.
When using LLMs for scientific research, we need some level of reliability, objectivity, and transparency in the process. To do so, we need to turn that process of prompt revision into an objectively defined and replicable set of steps. To overcome the issues of personal biases, subjectivity, and inconsistencies, we need to do the following.
  1. (1)
    Ensure that at least two qualified researchers are involved in the whole process.
  2. (2)
    Before focusing on revising the prompt to produce desirable outcomes, clearly and objectively articulate what makes a desirable outcome. This will involve establishing a reliable and if possible, community-validated metric for assessment.
  3. (3)
    Discuss individual differences and biases with a goal of rooting them out and creating assessments and a set of instructions that can be understood and carried out by any researcher with a reasonable set of skills to be able to replicate the experiments and achieve similar results.
While a few recent works have shown promise in addressing these desiderata, they still lack enough rigor to meet the requirements of transparency, objectivity, and replicability. For instance, Shah et al. (Shah et al., 2023b) use a verifying method for prompt generation, but their approach lack verification of criteria as well as replicability for different datasets.
3. Methodology
In this section we describe a new methodology for making prompt engineering efforts to a scientific process that ensures reliability and generalizability. The previous section described how a typical prompt engineering process works; essentially, it involves revising the prompt to an LLM in order to fit the desired response and/or the downstream application. We identified that one of the problems with this endeavor is that it is quite ad-hoc, unreliable, and not generalizable. Even with enough documentation of how the final prompt was generated, one could not ensure that the same process will yield same quality output in a different situation – be it with a different LLM or the same LLM in a different time. To overcome these problems, we describe an approach that first iterates over the criteria for assessing LLM’s response and then iterates over the prompt in a systematic way using human in the loop. Our method is inspired by a typical process of qualitative coding.
3.1. Qualitative coding with multiple assessors
The problem we are trying to address here is that of construction of a prompt or a set of instructions that produces desired outcomes for unseen data. This is similar to building a codebook in qualitative research. Take for instance, work described in (Mitsui et al., 2016). Here, the authors were attempting to label user behaviors (data) with user intents (discrete labels). For this, they needed to identify a finite (and preferably a small) set of labels, and their desired descriptions. This is called a codebook.
They began their process by having an initial codebook that was inspired by prior work and literature. But now they need to fine-tune and validate that codebook for the given application. For this, they followed the qualitative coding process depicted in Figures 2 and 3.
Figure 2. Training portion of qualitative coding.
Figure 3. Testing portion of qualitative coding.
As shown, for a typical qualitative coding process two researchers use the initial codebook and a small set of data to independently provide labels during the training portion. The process continues until there is good enough agreement or inter-coder reliability (ICR). Once that point is reached, the codebook is considered to be ready for use. It can now be used by the same or other researchers to independently and non-overlapping label unseen data during the testing portion.
It is important to note that the value of this iterative process with human-in-the-loop. During this process, it is not just the codebook that is getting revised, but also the understanding in those two researchers. Through their discussions of disagreements, they are able to think through the task, the data, and the best ways to describe that data. This enhanced understanding impacts the codebook and vice versa. In other words, this process is used not only to produce, but also to learn and discover for the researchers. But this is not a simple form of brainstorming. The researchers are discussing their differences and charting their way forward bound by a rigorous set of criteria. For instance, in this example, the researchers were aiming to produce a codebook with categories that were clear, concise, comprehensive, and reliable. They had specific definitions of these criteria, which they used during their deliberations.
In short, such a process of qualitative coding ensures that individual subjectivity is dissolved to the extend possible while producing a coding scheme that is robust, verifiable, and explainable. Next, we show how this can be used for constructing a method for prompt generation and verification.
3.2. Codebook building as a way to construct prompts
We will now describe a new method for prompt generation inspired by the qualitative codebook building process described above. Once again, our objective is to generate a prompt for a given LLM in order to achieve desired response. In addition, we need this prompt to be systematically created with enough generalizability and explainability to be useful for same or similar experiments by other researchers, by other LLMs, and in different times. We will describe the abstract process in this section and its operationalization in the next section. To begin, we list two primary desiderata:
  1. (1)
    The desired response from the LLM must be verified along a set of dimensions and for large enough sample size by at least two humans.
  2. (2)
    The prompt to generate must consistently generate the desired responses and must be justifiable, explainable, and verifiable.
To operationalize these two criteria, we propose a multi-phase process, working backward in the pipeline of prompt to application as shown in Figure 4. There are four phases to go through in this method, each explained below.
Figure 4. Proposed methodology with multi-phase approach to response and prompt validations.
Phase 1: Set up the initial pipeline.
This can be done using anything that is reasonable. For instance, the initial prompt could be a simple and direct instruction to generate the desired response.
Phase 2: Identify appropriate criteria to evaluate the responses.
Before we can tweak the prompt, it is important to establish that we know how to evaluate the response. For this, a set of appropriate criteria could be obtained from prior work or literature. But these criteria still need to be tested and validated for the given application. Moreover, eventually we will need human assessors to assess if the LLM is generating desired responses in order to revise the prompt. For these reasons, we need to execute our first human-in-the-loop assessment. To do so, follow the steps below.
  1. (1)
    Run the existing pipeline to generate a reasonable number of responses. Give each assessors these responses, the existing set of criteria (codebook), and sufficient details of the application. The assessors should review these things before starting their assessment. This can be supervised by the lead researcher or an expert.
  2. (2)
    Each assessor applies the criteria on the given responses independently.
  3. (3)
    Compare the assessors’ outcomes and compute the level of agreement. This could be using as simple as measuring the % times when they agree, or using inter-coder reliability (ICR) using an appropriate method such as Cohen’s kappa (Warrens, 2015) or Krippendorfff’s alpha (Krippendorff, 2011).
  4. (4)
    If the amount of agreement is not sufficient (this could vary from application to application), bring the assessors together to discuss their disagreement. This is a very important step as it offers an opportunity to not only remove individual biases and subjectivity, but also influences researchers’ collective understanding of the task and the criteria for evaluation. The discussion could result in resolution of conflict, but also changes in the codebook. These changes could be about the number or the description of the criteria for assessment. Now, a new set of responses are generated for assessment using the revised codebook and the above process is repeated.
  5. (5)
    Once a sufficient level of agreement is achieved, the codebook for assessing the responses is finalized. At this point the assessors may still choose to make minor changes in the codebook to enhance its readability with the idea that anyone who did not participate in this process of developing it can still understand and use it very much how those assessors would.
Phase 3: Iteratively develop the prompt.
Continuing to work backward, now we came to the next phase of codebook development, which happens to be the first phase of the pipeline depicted in Figure 4. The objective here is to see if the LLM, with the current version of the prompt, is meeting our criteria for the responses. If not, we need to figure out what changes can be made to the prompt.
At this point, we know what makes a good response from our LLM. More importantly, we have now trained at least two researchers in reasonably objectively, consistently, and independently evaluating such responses. They could more effectively assess what instructions could help produce desired responses. Follow the steps below to iteratively develop the prompt.
  1. (1)
    Run a reasonable number of data inputs using the current prompt.
  2. (2)
    Give the generated responses to the human assessors, ideally the same ones who were involved in the previous phase. However, this is not a requirement. These assessors should use the codebook produced by the previous phase to assess the responses.
  3. (3)
    Have them assess if the generated responses meet the criteria for the application in question.
  4. (4)
    Compare the assessments by different assessors to find their level of agreement. If that agreement is not sufficient, have the assessors discuss their disagreements, possibly under the supervision of a senior researcher or an expert, and revise the codebook as necessary. Here, the codebook is the prompt fed to the LLM. Repeat the above process.
  5. (5)
    Once there is enough agreement among the assessors that the prompt meets the objective of generating desired responses more than some threshold, this process is done. The assessors or the supervising researcher may still make minor adjustments to the prompt for improved readability, interpretability, and generalization.
Phase 4: Validate the whole pipeline.
As an optional and final validation phase, run through the whole pipeline using a portion of the test data and evaluate random samples of the responses to ensure the whole process still yields quality results that can be independently and objectively validated. For this, it is ideal to use a different set of assessors and have them independently label that same set of randomly sampled responses. Compute their ICR on this sample to see if (1) there is a good enough agreement among them on labeling; and (2) the generated responses are meeting the desired criteria. If either of these objectives are not met, take appropriate actions for correction, based on phases 2 and 3.
As one can see, there are multiple checkpoints and validations in this method that ensures removal of ad-hocness, subjectivity, and opaqueness as much as possible. A proper documentation of its execution will yield a more informative, meaningful, and scientific communication that can help other researchers to trust the outputs as well as test and build such pipelines themselves.
4. Experiments
As a demonstration of the methodology explained in the previous section, we will now walk through its execution for a specific research application. The problem to address in this research is that of auditing an LLM. While there have been many recent efforts in this area (Mökander et al., 2023; Rastogi et al., 2023), we still lack systematic and scientifically rigorous approaches. We attempted to address this using a simple, but effective method – prompting the same question differently to the given LLM to see how differently it responds. A detailed discussion of the auditing problem for an LLM is beyond the scope for the present article. Instead, our focus here is on constructing the set of probes that could be used for that auditing task.
Let us start by identifying different components for the pipeline in Figure 4.
  • Data: Original set of questions that we could use to probe an LLM.
  • Responses: Different prompts generated using the original question. These prompts can be then given to the LLM being investigated.
  • Criteria: We need to make sure that the responses generated are still representing the original question, and yet different enough to be meaningful for probing an LLM. In other words, our main criteria here are relevance (to the original question) and diversity (among the generated responses/prompts).
  • Application: We will use the responses that meet our criteria of relevance and diversity as prompts to the LLM being investigated. The idea here is if the original question calls for factual responses and if the LLM being audited gives meaningfully different answers with that question asked differently or by different people/persona, then the said LLM need further probing as there may be issues of bias or hallucination.
To operationalize the method proposed in Section 3.2, we started with setting up the initial version of the pipeline. The initial prompt was set as “Generate five different questions for the following question that represent the same meaning, but are different.” for the LLM, we used Mistral 7-B and the data for original questions came from TruthfulQA dataset (Lin et al., 2021). Given the application – generating different versions of the same question for auditing an LLM – we set the criteria for responses generated by Mistral as relevance and diversity. With this, phase 1 was over and we moved to phase 2.
For phase 2, we used two researchers with a good understanding of the application and qualitative assessment. They were tasked with iteratively developing and finalizing the criteria for assessing the generated responses (different versions of the original question). The assessors started with a 7-point Likert scale for assessing relevance and a binary scale for assessing diversity. 10 original questions were sent through the LLM with the initial prompt to generate 10 sets of responses. We decided to limit each set to contain five variations. The two assessors labeled these 10 sets separately. They marked for ‘relevance’ for 50 questions and marked for ‘diversity’ for 10 sets.
Once done, we measured their agreement, which turned out to be very low. The problem seemed to stem from 7-levels of relevance scale and the two researchers’ understanding of what each level meant. This resolution for the scale did not seem necessary. We, therefore, decided to change the scale for relevance to three levels: low-medium-high. Another round of coding on 10 different input data was done. Once again, we had relevance assessment for 50 data points and diversity assessment for 10 data points.
We used Cohen’s kappa for measuring ICR, and found that the values for this measure for relevance was 0.4634 and for diversity was 0.6154. These are considered to be moderate levels of agreements (Ritchie et al., 2003). After a discussion between the two researchers, the definitions of relevance and diversity were slightly revised and the coding process was repeated with a new set of data. After one more iteration, we achieved 0.6418 and 0.8276 for Cohen’s kappa values for relevance and diversity, respectively. These are considered to be substantial levels of agreement (Ritchie et al., 2003). At this point, we decided to finish phase 2. The primary outcome was the codebook that contained detailed instructions about how to evaluate relevance and diversity in this particular context. An important byproduct was the enhanced understanding that the two researchers now possessed for assessing the outputs of the LLM in a reasonably objective and consistent manner.
Figure 5. Final prompt template.
We now proceeded with phase 3. The criteria for evaluating the outputs of the LLM was set. We needed to now focus on revising the prompt. Specifically, we wanted to have a prompt that consistently generates outputs (a set of questions) that could achieve high enough relevance and diversity for their use in a downstream application. We set the threshold for this to be 75%. Similar to the previous phase, the same two researchers marked the generated responses for relevance and diversity using a new set of input data. The initial round of assessment showed that only around 50% of the outcomes met the goal of relevance (medium or high) and diversity. To change this, the original prompt was modified to include more details about the application and the criteria. Then this process was repeated. After two more iterations, the LLM was generating outcomes that were relevant and diverse for at least 80% of the cases. At this point, the process for phase 3 was complete.
For phase 4, a senior researcher tested the whole pipeline with the final versions of the prompt and the criteria with a few rounds of random samplings. They made a few minor adjustments to the prompt for improved readability and generalizability. The final prompt is shown in Figure 5. At this point, we can be assured that this prompt will lead to reasonably consistent, reliable, and high quality outputs. The prompt can also be easily modified to suit small variations in the downstream applications without having to repeat the whole method. The documentations from the discussions and design choices during phases 2 and 3 could be shared, much like source code for an application, to aid in explainability and extensibility of this pipeline.
5. Conclusion
LLMs could simply be stochastic string generation tools (Bender et al., 2021) that are effective at predicting next tokens. But for many researchers, they are proving to be useful in many research problems that in general involve labeling or classifying input data based on some analysis or reasoning. However, using LLMs blindly for such research tasks just because their outputs seem reasonable can be dangerous. It can perpetual biases (Shah and Bender, 2022), create fabricated or hallucinated responses (Rawte et al., 2023), and provide unverifiable results (Stechly et al., 2023). All of these are harmful for scientific progress.
Creatively or iteratively designing prompts to produce desirabile outcomes, what is often referred to as prompt engineering, does not address these concerns. In fact, prompt engineering may even worsen the problems by focusing too much on bending the process to generating the desired outcomes than to developing a verifiable, validated, and replicable process. In this article, we showed how prompt engineering can be turned into prompt science by introducing a multi-phase process with human-in-the-loop. Specifically, we divided the prompt to application pipeline into two major phases that provide two levels of verification: one for the criteria for evaluating the responses generated by the LLM, and the other for the prompt for generating the responses. Through a method inspired by qualitative coding for codebook development, we showed how to have scientific rigor in developing a reliable prompt and getting trustworthy response for a downstream application. This method was demonstrated using a case study that used an LLM for generating questions for auditing another LLM.
We believe there are three main reasons this method provides scientific rigor.
  1. (1)
    It provides a systematic and scientifically verifiable way of producing prompt templates, while removing individual subjectivity and biases.
  2. (2)
    Through the involvement of multiple researchers documenting their deliberations and decisions, it fosters openness and replicability. This documentation can and should be made available, much like the code for an open source tool, allowing other researchers to validate, replicate, and extend the work.
  3. (3)
    In addition to producing a prompt template that can generate desired outcomes reliably and consistently, the process also adds to our knowledge about the problem at hand (application), how to best assess what we need for that problem, and how we could consistently and objectively analyze given data.
Such a rigorous process comes at a cost. Comparing the typical process of prompt engineering (Figure 1) to the proposed method of prompt science (Figure 4), there is at least a 3-fold cost increase. Guided deliberations and detailed documentation can also add to the cost. Finally, the optional phase 4 calls for a lightweight, but more continual cost to the whole pipeline to maintain its quality and validity. We can compare this with an assembly line that is primarily run through automated processes with some human supervision. Depending on the risk of bad quality and the cost for manual quality check, one could decide how often and how rigorously humans should be involved. Considering that the proposed method not only leads to a outputs with higher quality, consistency, and reliability, but also an increased understanding of the task and the data, we believe the added cost is justifiable. The broader scientific community could also benefit from exercising and perhaps insisting on such a rigor so we could leverage LLMs ethically and responsibly in our research.
Acknowledgements.
The experiments described in Section 4 are results of hard work by Maryam Amirizaniani, Adrian Lavergne, Elizabeth Okada, and Jihan Yao. We are also thankful to Aman Chadha and Tanya Roosta for their valuable comments and guidance during this process.